OpenStudy (anonymous):

A sequence is defined recursively using the equation f(n + 1) = f(n) – 8. If f(1) = 100, what is f(6)?

OpenStudy (p0sitr0n):

make a table: f(1)=100 f(2)=100-8=92 now you observe that f(i)=100-(i-1)*8 so now you take i=6, f(6)=100-(6*8)=100-48 what you basically do is convert a recursion into a normal sequence that does not depend on previous terms.

OpenStudy (p0sitr0n):

my bad, f(6)=100-(5*8)=100-40

OpenStudy (anonymous):

I got it thanks alot.

OpenStudy (p0sitr0n):

np